# Learning to Think: Information-Theoretic Reinforcement Fine-Tuning for LLMs

### v2

@@ C.2 More Comparison
Firstly, previous methods [43, 16, 53, 9, 49] mainly optimize via outcome rewards, which drives models to over-extend reasoning chains and waste test-time compute. In contrast, our dense process rewards immediately quantify each episode’s contribution to performance, enabling the model to learn when to stop reasoning and thus achieve equal or better accuracy with a minimal token budget.
Secondly, some concurrent works [57, 1, 33, 36, 20] propose task-specific process rewards, heuristic scorers, and length penalties to reduce the length of the reasoning chains, helping save test-time compute. However, these methods require costly manual labeling and do not transfer across tasks since they rely on task-specific settings, and there is no one-size-fits-all solution. We instead measure the parameter-update signal inspired by information theory, i.e., the intrinsic change in the model’s weights after each optimization, as a task-agnostic proxy for learning progress. This internal metric requires no additional annotations or retraining and applies uniformly across diverse reasoning tasks.
Thirdly, existing reward-based updates to optimize test-time compute [36, 57, 1, 40, 64] may embed noise or task-specific artifacts into model weights, leading to overfitting and drastic performance drops under slight input shifts. In contrast, we introduce a parameter compression penalty that quantifies and negatively rewards the absorption of redundant information at each update. Only updates yielding true information gain are amplified; small or harmful directions are suppressed, ensuring stability under noisy, ambiguous, or adversarial inputs. Moreover, by recasting LLM fine-tuning as a episodic RL problem, combining a task distribution with dense per-step feedback, we enable the policy to maintain robust performance from simple arithmetic to complex proofs, without redesigning rewards or retuning hyperparameters for each new task.
@@
C.3 Broader Impacts and Limitations
@@
Broader Impacts. This work explores how to simultaneously maximize inference effectiveness and efficiency across tasks of varying complexity to meet real-world demands. It reformulates LLM reasoning through a episodic reinforcement-learning and information-theoretic lens, and introduces a general dense process reward to track reasoning progress, enabling optimal performance under a minimal token budget. Extensive theoretical and empirical analyses validate its effectiveness and robustness. This work also opens up exciting new avenues for future research.
Limitations. This study evaluates general LLM reasoning tasks, such as mathematical proofs and code generation, which are commonly used to verify the reasoning capability of LLMs; some newly proposed benchmarks, such as web design, were not used. Our current experiments use mainly the open-source DeepSeek base models, including the scales of 1.5B, 3B, and 7B, while the scale of 72B and even above 100B is not used due to resource limitations and not being open source. We will investigate additional case studies and more base models to extend this work in the future.
Appendix D Benchmark Datasets
@@
  • AIME24-25 comprises 30 questions from the 2024 and 2025 American Invitational Mathematics Examination, with 15 fill-in-the-blank questions per exam. These questions are more difficult than AMC, spanning number theory, combinatorics, geometry, and algebra.
  • AMC10/12 consists of 25 multiple-choice questions each for the AMC10 (up to 10th grade) and AMC12 (up to 12th grade). Each competition consists of 25 multiple-choice questions, totaling 975 questions across 39 tests. Questions progress from basic algebra and geometry to introductory probability and counting, covering various tasks for LLM reasoning evaluation.
  • MATH500 is a 500-question subset randomly sampled from MATH, covering seven topics—prealgebra, algebra, number theory, geometry, intermediate algebra, precalculus, etc. Each question includes a step-by-step solution and a difficulty label from 1 to 5, enabling evaluation of an LLM’s mathematical question-solving across diverse domains.
  • MinervaMATH comprises 12,500 high-school-level contest questions. Each includes detailed solution steps and spans prealgebra through precalculus.
  • Omni-MATH is an Olympiad-level benchmark of 4,428 competition questions across 33 subdomains (e.g., number theory, combinatorics, geometry, algebra), stratified into over 10 difficulty levels (divided into 4 tiers following [5]).
@@
Appendix E Implementation Details
The implementation of our approach is built upon the openRL codebase and Llama-Factory codebase, with custom modifications to the loss function, i.e., replacing the reward function to our proposed reward mentioned in Subsection 4.2. For model training, we directly load the base models from Hugging Face, including DeepScaleR-1.5B-Preview, DeepSeek-R1-Distill-Qwen-1.5B, DeepSeekR1-Distill-Qwen-7B, and Qwen2-7B-Instruct. For different reasoning tasks, we introduce the experimental settings in the corresponding sections of Section 5 and Appendix F. Unless otherwise specified, we follow the protocol of each benchmark and record the maj@4 results across different models. The training configuration is: the learning rate is set to $1.0\times 10^{-6}$, with a cosine learning rate scheduler, and a warm-up ratio of 0.1. We use a batch size of 256, with a maximum prompt length of 4,096 tokens and a maximum completion length of 16,384 tokens. The model is trained for 1 epoch, up to 10 epochs. Additionally, we set the ‘use_vllm’ flag to True to enable vLLM acceleration, with a GPU memory utilization of 0.8. We also utilize mixed precision training with BF16 enabled. The parameters for compression penalty approximation are handled by a single-layer MLP, with a Fisher information matrix damping factor set to $10^{-5}$. The regularization hyperparameters $\alpha$ and $\beta$ are set to 0.8 and 0.6 according to grid search results, respectively. Also, $\alpha$ can be set to 1 for simplicity, with the performance drop less than 1%. More evaluation of implementation is provided in Appendix F. The entire training is conducted on A100 GPU clusters, ensuring scalability and high computational efficiency. The main code is provided in supplementary materials.
Appendix F Additional Experiments and Full Results
@@
DeepScaleR-1.5B-Preview 42.8 36.7 83.0 85.2 24.6 54.5
  +outcome-reward RL (GRPO) 44.5 (+1.7) 39.3 (+2.6) 81.5 (-1.5) 84.9 (-0.3) 24.7 (+0.1) 55.0 (+0.5)
@@
  +outcome-reward RL (GRPO) 29.8 (+1.1) 27.3 (+1.3) 70.5 (+0.6) 80.3 (+0.2) 22.1 (+2.3) 46.0 (+1.1)
@@
DeepSeek-R1-Distill-Qwen-7B 55.5 50.2 80.5 83.3 42.1 62.3
  +outcome-reward RL (GRPO) 56.9 (+1.4) 51.7 (+1.5) 81.0 (+0.5) 83.6 (+0.3) 43.5 (+1.4) 63.3 (+1.0)
  +length penalty 53.8 (-1.7) 46.9 (-3.3) 77.0 (-3.5) 80.8 (-2.5) 39.5 (-2.6) 59.6 (-2.7)
  +MRT-Reproduct 57.0 (+1.5) 52.4 (+2.2) 82.8 (+2.3) 84.0 (+0.7) 44.3 (+2.2) 64.1 (+1.8)
  +Ours 58.4 (+2.9) 53.6 (+3.4) 83.0 (+2.5) 85.5 (+2.2) 45.0 (+2.9) 65.1 (+2.8)
@@
Figure 7: Efficiency comparison on DeepScaleR-1.5B-Preview (a-d), DeepSeek-R1-Distill-Qwen-1.5B (e-h), and DeepSeek-R1-Distill-Qwen-7B (i-l). We compute the token budget required for each benchmark and treat the budget of the base model w/o fine-tuning as reference (1×1$\times$1 ×).
@@
F.1 More Details and Results of the Motivating Experiments
@@
F.2 Full Results of Effectiveness and Efficiency Analysis
To evaluate the proposed L2T, we conduct experiments on various benchmarks, including mathematical and code-related tasks, and across base models of different scales. In the main text, we report the performance of L2T on mathematical reasoning tasks of varying complexity. Table 1, Figure 2, and Figure 3 show that, compared to prior outcome-reward or process-reward methods, L2T delivers superior reasoning with less test-time compute. In this subsection, we assess its performance on more based models and more reasoning tasks, e.g., code generation tasks. Notably, mathematical reasoning and code generation serve as classic benchmarks for testing an LLM’s complex reasoning ability [43, 16, 32]. First, we report performance across additional model scales. The results for inference accuracy and compute efficiency are shown in Table 2 and Figure 7. We observe the same conclusion: L2T achieves state-of-the-art performance, attaining the highest inference accuracy with the smallest token budget. These findings demonstrate the broad effectiveness of our approach across models of varying scales. Secondly, we provide the performance of the proposed framework on the code generation task, and we evaluate the improvement of L2T on the LLM reasoning performance on the code generation task. Specifically, we run the GRPO/L2T fine-tuning pipeline on Qwen2-7B-Instruct and evaluate it using the standard HumanEval protocol. We set the temperature to 0.6 and top-p to 0.95 and generate 64 solutions per question, with a 1s timeout per attempt. We report the proportion of questions that pass all unit tests at least once. From the results in Figure 8, we can observe that compared with the outcome-reward-based RL method, L2T achieves better reasoning performance with less token budget. This proves the superiority of our proposed L2T and dense process rewards.
F.3 Full Results of Ablation Studies

### v3

@@ D.3 More Comparison
Firstly, previous methods [43, 16, 53, 9, 49] mainly optimize via outcome rewards, which drives models to over‑extend reasoning chains and waste test‑time compute. In contrast, our dense process rewards immediately quantify each episode’s contribution to performance, enabling the model to learn when to stop reasoning and thus achieve equal or better accuracy with a minimal token budget.
Secondly, some concurrent works [57, 1, 33, 36, 20] propose task‑specific process rewards, heuristic scorers, and length penalties to reduce the length of the reasoning chains, helping save test-time compute. However, these methods require costly manual labeling and do not transfer across tasks since they rely on task-specific settings, and there is no one-size-fits-all solution. We instead measure the parameter‑update signal inspired by information theory, i.e., the intrinsic change in the model’s weights after each optimization, as a task‑agnostic proxy for learning progress. This internal metric requires no additional annotations or retraining and applies uniformly across diverse reasoning tasks.
Thirdly, existing reward‑based updates to optimize test-time compute [36, 57, 1, 40, 64] may embed noise or task‑specific artifacts into model weights, leading to overfitting and drastic performance drops under slight input shifts. In contrast, we introduce a parameter compression penalty that quantifies and negatively rewards the absorption of redundant information at each update. Only updates yielding true information gain are amplified; small or harmful directions are suppressed, ensuring stability under noisy, ambiguous, or adversarial inputs. Moreover, by recasting LLM fine‑tuning as a episodic RL problem, combining a task distribution with dense per‑step feedback, we enable the policy to maintain robust performance from simple arithmetic to complex proofs, without redesigning rewards or retuning hyperparameters for each new task.
@@
D.4 Broader Impacts and Limitations
@@
Broader Impacts. This work explores how to simultaneously maximize inference effectiveness and efficiency across tasks of varying complexity to meet real‑world demands. It reformulates LLM reasoning through a episodic reinforcement‑learning and information‑theoretic lens, and introduces a general dense process reward to track reasoning progress, enabling optimal performance under a minimal token budget. Extensive theoretical and empirical analyses validate its effectiveness and robustness. This work also opens up exciting new avenues for future research, e.g., provides a way for more explicit and automated dynamic budget allocation in the future, especially in scenarios sensitive to token costs (e.g., mobile deployment or real-time QA).
Limitations. This study evaluates general LLM reasoning tasks, such as mathematical proofs and code generation, which are commonly used to verify the reasoning capability of LLMs; some newly proposed benchmarks, such as web design, were not used. Our current experiments use mainly the open‑source DeepSeek base models, including the scales of 1.5B, 3B, and 7B, while the scale of 72B and even above 100B is not used due to resource limitations and not being open source. We will investigate additional case studies and more base models to extend this work in the future.
Appendix E Benchmark Datasets
@@
  • AIME24-25 comprises 30 questions from the 2024 and 2025 American Invitational Mathematics Examination, with 15 fill‑in‑the‑blank questions per exam. These questions are more difficult than AMC, spanning number theory, combinatorics, geometry, and algebra.
  • AMC10/12 consists of 25 multiple‑choice questions each for the AMC10 (up to 10th grade) and AMC12 (up to 12th grade). Each competition consists of 25 multiple-choice questions, totaling 975 questions across 39 tests. Questions progress from basic algebra and geometry to introductory probability and counting, covering various tasks for LLM reasoning evaluation.
  • MATH500 is a 500‑question subset randomly sampled from MATH, covering seven topics—prealgebra, algebra, number theory, geometry, intermediate algebra, precalculus, etc. Each question includes a step‑by‑step solution and a difficulty label from 1 to 5, enabling evaluation of an LLM’s mathematical question‑solving across diverse domains.
  • MinervaMATH comprises 12,500 high‑school-level contest questions. Each includes detailed solution steps and spans prealgebra through precalculus.
  • Omni‑MATH is an Olympiad‑level benchmark of 4,428 competition questions across 33 subdomains (e.g., number theory, combinatorics, geometry, algebra), stratified into over 10 difficulty levels (divided into 4 tiers following [5]).
@@
Appendix F Implementation Details
For model training, we directly load the base models from Hugging Face, including DeepScaleR-1.5B-Preview, DeepSeek-R1-Distill-Qwen-1.5B, DeepSeekR1-Distill-Qwen-7B, and Qwen2-7B-Instruct. For different reasoning tasks, we introduce the experimental settings in the corresponding sections of Section 5 and Appendix G. Unless otherwise specified, we follow the protocol of each benchmark and record the maj@4 results across different models. The training configuration is: the learning rate is set to $1.0\times 10^{-6}$, with a cosine learning rate scheduler, and a warm-up ratio of 0.1. We use a batch size of 256, with a maximum prompt length of 4,096 tokens and a maximum completion length of 16,384 tokens. The model is trained for 1 epoch, up to 10 epochs. Additionally, we set the ‘use_vllm’ flag to True to enable vLLM acceleration, with a GPU memory utilization of 0.8. We also utilize mixed precision training with BF16 enabled. The parameters for compression penalty approximation are handled by a single-layer MLP, with a Fisher information matrix damping factor set to $10^{-5}$. The regularization hyperparameters $\alpha$ and $\beta$ are set to 0.8 and 0.6 according to grid search results, respectively. Also, $\alpha$ can be set to 1 for simplicity, with the performance drop less than 1%. More evaluation of implementation is provided in Appendix G, e.g., parameter sensitivity, prompt configuration, etc. The entire training is conducted on A100 GPU clusters, ensuring scalability and high computational efficiency.
Appendix G Additional Experiments and Full Results
@@
DeepScaleR‑1.5B‑Preview 42.8 36.7 83.0 85.2 24.6 54.5
+outcome‑reward RL (GRPO) 44.5 (+1.7) 39.3 (+2.6) 81.5 (-1.5) 84.9 (-0.3) 24.7 (+0.1) 55.0 (+0.5)
@@
+outcome‑reward RL (GRPO) 29.8 (+1.1) 27.3 (+1.3) 70.5 (+0.6) 80.3 (+0.2) 22.1 (+2.3) 46.0 (+1.1)
@@
DeepSeek-R1-Distill-Qwen-7B 55.5 50.2 85.1 87.4 42.1 64.1
+outcome-reward RL (GRPO) 56.9 (+1.4) 51.7 (+1.5) 85.5 (+0.4) 87.7 (+0.3) 43.5 (+1.4) 65.1 (+1.0)
+length penalty 53.8 (-1.7) 46.9 (-3.3) 81.2 (-3.9) 83.7 (-3.7) 39.5 (-2.6) 61.0 (-3.1)
+MRT-Reproduct 57.0 (+1.5) 52.4 (+2.2) 86.0 (+0.9) 88.4 (+1.0) 44.3 (+2.2) 65.6 (+1.5)
+Ours 58.4 (+2.9) 53.6 (+3.4) 87.5 (+2.4) 89.2 (+1.8) 45.0 (+2.9) 66.8 (+2.7)
@@
Figure 7: Efficiency comparison on DeepScaleR-1.5B-Preview (a-d), DeepSeek-R1-Distill-Qwen-1.5B (e-h), and DeepSeek-R1-Distill-Qwen-7B (i-l). We compute the token budget required for each benchmark and treat the budget of the base model w/o fine-tuning as reference ($1\times$).
@@
(a) Simple Arithmetic
(b) Fraction Addition
(c) Geometry
(d) Geometry
Figure 10: Examples of qualitative analysis about $\alpha$.
G.1 More Details and Results of the Motivating Experiments
@@
G.2 Full Results of Effectiveness and Efficiency Analysis
To evaluate the proposed L2T, we conduct experiments on various benchmarks, including mathematical and code-related tasks, and across base models of different scales. In the main text, we report the performance of L2T on mathematical reasoning tasks of varying complexity. Table 1, Figure 2, and Figure 3 show that, compared to prior outcome-reward or process-reward methods, L2T delivers superior reasoning with less test-time compute. In this subsection, we assess its performance on more base models and more reasoning tasks, e.g., code generation tasks. Notably, mathematical reasoning and code generation serve as classic benchmarks for testing an LLM’s complex reasoning ability [43, 16, 32]. First, we report performance across additional model scales. The results for inference accuracy and compute efficiency are shown in Table 2 and Figure 7. We observe the same conclusion: L2T achieves state-of-the-art performance, attaining the highest inference accuracy with the smallest token budget. These findings demonstrate the broad effectiveness of our approach across models of varying scales. Secondly, we provide the performance of the proposed framework on the code generation task, and we evaluate the improvement of L2T on the LLM reasoning performance on the code generation task. Specifically, we run the GRPO/L2T fine-tuning pipeline on Qwen2-7B-Instruct and evaluate it using the standard HumanEval protocol. We set the temperature to 0.6 and top-p to 0.95 and generate 64 solutions per question, with a 1s timeout per attempt. We report the proportion of questions that pass all unit tests at least once. From the results in Figure 8, we can observe that compared with the outcome-reward-based RL method, L2T achieves better reasoning performance with less token budget. This proves the superiority of our proposed L2T and dense process rewards.
To further validate the versatility of our approach, we conduct two complementary studies. First, since L2T improves reasoning efficiency by adaptively allocating token budgets, we examine whether additional test-time search (e.g., MCTS) provides further gains. We apply MCTS to models fine-tuned with GRPO and L2T, and evaluate pass@1 accuracy on AIME and MinervaMATE. The results in Table 4 show that GRPO benefits noticeably from MCTS, while L2T already achieves strong reasoning performance and gains only marginally, confirming that L2T effectively reduces reliance on large-scale search. Second, as our dense process reward is defined in a task-agnostic form, we also apply it to inference-only methods as a reranking signal. We evaluate DeepSeek-R1-Distill-Qwen-7B under a best-of-16 setting on MinervaMATH, comparing log-likelihood reranking, PRM-based reranking, and L2T reranking. The results in Table 4 demonstrate that L2T achieves the highest accuracy, outperforming both likelihood-based and PRM baselines. Together, these findings indicate that L2T not only enables efficient reasoning during fine-tuning but also serves as an effective scoring mechanism in inference-time pipelines.
Method AIME 2024 MinervaMATE
GRPO 44.5 24.7
GRPO+MCTS 46.8 25.3
L2T 48.5 26.5
L2T+MCTS 48.9 26.7
Table 3: Effect of combining with MCTS.
Method MinervaMATH
Best-of-16 + log-likelihood 42.6
Best-of-16 + Qwen2.5-Math-PRM 43.3
Best-of-16 + Skywork-PRM 43.0
Best-of-16 + L2T rerank 43.9
Table 4: Applying to inference-only reranking.
G.3 Full Results of Ablation Studies
@@
G.4 Visualization
G.4.1 Qualitative Analysis of $\alpha$
In this subsection, we construct qualitative experiments to illustrate how to eliminate redundant inferences at low and high alpha values. The coefficient $\alpha$ denotes the weight of our proposed process reward (Eq.5). When $\alpha$ is small, the model relies more on the outcome reward, which provides high rewards only when the correct answer is found. Without guidance and given that correct answers are sparse in the output space, the model may consume a large number of tokens in exploration, reducing efficiency. In contrast, when $\alpha$ is large, the model is driven by the process reward, which assigns high rewards only if the current reasoning step has a high contribution to the accuracy of the answer, i.e., the correct answer is reached within a short token sequence. This encourages the model to generate informative tokens at each step, thereby improving efficiency. The qualitative results are shown in Figure 10, which demonstrate the above analyses. Take the Tier-1 Omni-MATH problem “What is 12 + 5?” as an example, the qualitative analysis shows that with $\alpha=0.9$, the model may answer within 2–3 steps, whereas with $\alpha=0.2$, it may take more than 7 steps.
G.4.2 Prompt Configuration for Episode Segmentation
For episode segmentation, we automatically segment the chains by designing specific prompts.
Taking mathematical reasoning tasks as an example, to segment the reasoning chain into fixed episodes (e.g., “segment up to 30”), we add the following instruction to the prompt file:
For adaptive segmentation (e.g., “break the reasoning chains into $K$ segments”), we add:
G.4.3 Visualization of Episodes and Dense Reward
In practice, episode segmentation is guided by semantic completeness and logical separability. It is automatically handled by the model during generation and sampling via prompt instructions, rather than based on a fixed token count. As a result, episode lengths are variable. For example, each episode may correspond to a distinct reasoning action, such as “defining variables”, “substituting into equations”, or “structural inference”. We provide a brief example below:
To accommodate tasks with varying reasoning granularity: for problems with long reasoning chains (e.g., mathematical proofs or code generation), a single episode may span multiple tokens; in contrast, for shorter or more fragmented tasks (e.g., factual questions), episodes may be automatically compressed into short logical units without forced segmentation. Below, we provide two examples. A question with a short reasoning chain: A question with (relatively) long reasoning chains: